Trace the curve 𝑦 = 8/4-x^2 , and state all the properties you use to trace it. 


1
Expert's answer
2022-05-16T16:14:03-0400

"y=\frac{8}{4-x^2}"

1. "4-x^2 \not =0 \Rightarrow x \not= \plusmn 2"

Domain:

"(\u2212\u221e;\u22122)\u222a(\u22122,2)\u222a(2,\u221e)"

2. y-y− intersection:

x=0, "y=\frac{8}{4-(0)^2}=2"

Point (0, 2)


x− intersection(s):

y=0, "0=\frac{8}{4-x^2}"

No solution


There are no xx -intersections.

3. "y(-x)=\frac{8}{4-(-x)^2}=\frac{8}{4-x^3}=y(x)"

The function is even on its domain. The garph is symmetric with respect to yy -axis.

4. "lim_{\n\u200bx\u2192\u22122^- }\n(\\frac{8}{4-x^2})\n =\u2212\u221e"

"lim_{\n\u200bx\u2192\u22122^+}\n(\\frac{8}{4-x^2})\n =\u221e"

"lim_{\n\u200bx\u21922^- }\n(\\frac{8}{4-x^2})\n =\u221e"

"lim_{\n\u200bx\u21922^+ }\n(\\frac{8}{4-x^2})\n =\u2212\u221e"

Vertical asymptotes: x=-2, x=2.x=−2,x=2.

"lim_{\n\u200bx\u2192\u2212\u221e}\n(\\frac{8}{4-x^2})\n =0"

"lim_{\n\u200bx\u2192\u221e}\n(\\frac{8}{4-x^2})\n =0"

Horizontal asymptote: y=0.y=0.

5. "y'°(\frac{8}{4-x^2})'=\frac{16}{(4-x^2)^2}"

Find the critical number(s):

"y'=0 \Rightarrow \frac{16x}{(4-x^2)^2}=0 \Rightarrow x=0, x\not= \plusmn 2"

If x<-2, y'<0, y decreases.

If -2<x<0, y'<0, y decreases.

If 0<x<2, y'>0, y increases.

If x>2, y'>0, y increases.

"y(0)=\frac{8}{4-(0)^2}=2"

The function yy has a local minimum with value of 2 at x=0.

6. "y"=(\frac{16x}{(4-x^2)^2})'=\frac{16(4-x^2)^2-2x(4-x^2)(-2x)}{(4-x^2)^4}=\frac{16(4-x^2+4x^2}{(4-x^2)^3}=\frac{16(4+3x^2)}{(4-x^2)^3}"

If x<-2, y'<0, y is concave down.

If "-2<x<2, y''>0, y" is concave up.

If x>2, y''<0, y is concave down.

7. Sketch the graph.